OpenStudy (anonymous):

4. A Honda with a dealer invoice price of $19,700 was retail priced at $23,000. (a) How much is the percent markup based on selling price? Round to the nearest hundredth percent. (b) How much is the markup percent if cost had been the case (use the equivalent markup formula)? Round to the nearest hundredth percent.

OpenStudy (anonymous):

Percent markup = (difference between the two prices)/original price

OpenStudy (anonymous):

(23,000 - 19,700)/19,700 = 3300/19700 = 16.75% approx

OpenStudy (anonymous):

Thank you.